Project

General

Profile

Physics Shift Report

Start Date: 14th July
End Date: 18th July
Start Run: 07236
End Run: 07237
Physics shifter: Francois Drielsma

Summary

No data was taken from the 14th to the evening of the 18th of July because:
  • The priority is placed on the magnet preparation activities.
  • TOF1 was moved out of the beam line for required Tracker repairs.
  • TOF2, KL and the EMR moved upstream on Tuesday to be against the end of the cooling channel.
On the evening of the 18th from 8PM to 12AM, two reference runs were taken
  • TOF1, TOF2, KL and the EMR back in the beam line
  • No decay solenoid
Run number Run type Mom. setting Trigger condition Nbr. of spills Int. part. tr.
7236 Pion + 300 MeV/c TOF1 1770 37914
7237 Pion + 300 MeV/c TOF1 272 5834

Run documentation

Elog

Entries for each shift and major issues/faults. No problems found.

Run Condition Summary Spreadsheet/CDB

  • Both runs successfully uploaded to the grid and the CDB with all the run conditions
  • Integrated number of triggers and spills consistent between Spreadsheet and CDB.
  • Run 7236/7237 : mismatched target depth and no beam loss recorded:

1.0V beamloss in the spreadsheet, nothing in the CDB
42.05mm depth in the spreadsheet, 63.95 and 64.55mm in the CDB, respectively

Analysis

  • No particular analysis requested by the measurement coordinator
  • Online reconstruction plots produced manually for the downstream PID detectors (check the cabling, integrity)
    • The EMR is fully covered and the reconstruction is working well. It is off-centre along x by a good 10cm, should be asserted...
    • KL cabling seems OK, no strong inconsistency between the readouts at both ends, beam profile centred around the middle of the detector.
    • TOF2 events suspiciously distributed, are some of the HV off? Need to investigate. UPDATE: We do have signals coming out of the PMTs (see slab hits plot at the bottom of the page) but they are not reconstructed into a space point, probably from a lack of calibration in those channels.

Reconstruction Quality

  • The physics devil reports a "Download Error" for all the aforementioned runs. No plots were generated and the data were not reconstructed by it.
  • The files have been properly uploaded to the grid

imageEMR_occupancy.png (68.2 KB) Drielsma, François, 19 July 2015 11:44

imageEMR_recon.png (47.1 KB) Drielsma, François, 19 July 2015 11:44

imageKL.png (17.7 KB) Drielsma, François, 19 July 2015 12:17

imagetof_SpacePoints_XY_tof2.png (28.9 KB) Drielsma, François, 20 July 2015 11:13

imagetof_nSpacePoints_vs_spill.png (23.8 KB) Drielsma, François, 20 July 2015 11:13

imagetof_pmtHits_Plane0PMT0.png (21.1 KB) Drielsma, François, 20 July 2015 11:24

imagetof_pmtHits_Plane1PMT0.png (20.4 KB) Drielsma, François, 20 July 2015 11:24